OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. sine of pi divided by two times cosine of pi divided by seven plus cosine of pi divided by two times sine of pi divided by seven.

OpenStudy (anonymous):

cos(pi/5) cos(pi/7) +sin(pi/5) sin(pi/7)

OpenStudy (anonymous):

\[\sin (\Pi/2)\cos(\Pi/7)+\cos (\Pi/2)\sin(\Pi/7)\]

OpenStudy (anonymous):

thats the equation

OpenStudy (anonymous):

OH ok

OpenStudy (dayakar):

may be the above answer is wrong plz verify

OpenStudy (anonymous):

what do you mean?

OpenStudy (phi):

I would use sin(a+b)= sin(a) cos(b)+ cos(a) sin(b) (but "backwards")

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

Georgia your done

OpenStudy (dayakar):

\[\frac{ \sin \pi }{ 2\cos \frac{ \pi }{ ? }} + \frac{ \cos \pi }{ 2 \sin \frac{ \pi }{ 7 } }\]

OpenStudy (anonymous):

No Hawaii we beat you 0 and 55

OpenStudy (anonymous):

i dont think youre doing the right thing @dayakar

OpenStudy (anonymous):

So shut up Hawaii

OpenStudy (dayakar):

ok, i;m sorry

OpenStudy (phi):

you should get \[ \sin\left(\frac{\pi}{2}+\frac{\pi}{7}\right) = \sin\left(\frac{9\pi}{14}\right) \]

OpenStudy (anonymous):

That's exactly what I got thank you @phi